karen is 3 times as old as Bill, and Bill is 2 years younger than Janice. Six years ago Karen was 2.5 times the age that Janice is now. How old is each person now?
Answer by mananth(16946)

karen is 3 times as old as Bill, and Bill is 2 years younger than Janice. Six years ago Karen was 2.5 times the age that Janice is now. How old is each person now?
...
Janice be x years now
Bill will be x-2 years
Karen is 3(x-2)
...
six years ago karen will be
3(x-2)-6 years
...
3(x-2)-6 = 2.5x
3x-6-6=2.5x
3x-2.5x=12
0.5x=12
/0.5
x= 24 janice's age now
..
bill = x-2 = 22
..
Karen = 3(x-2) = 3(48-2)
=66 years
...
m.ananth@hotmail.ca